# A Criterion for Perfectoid Purity and the Rationality of Thresholds

### v1

@@ 6. Examples
With the notation as in 6.1, let $f=x_{1}^{N}+\cdots+x_{N}^{N}$ and assume $p>N$. For each $e\geq 1$, take an integer $0\leq s_{e}\leq N-2$ satisfying $s_{e}+1\equiv p^{e}\pmod{N}$. Then the splitting–order sequence of $f$ is $\boldsymbol{s}(f)=(0,s_{1},s_{2},\ldots)$. In particular, $A/f$ is perfectoid pure.
@@
With the notation as in 6.1, assume that $f\in\mathfrak{m}^{[p]}$.
@@
    Then $A/f$ is not perfectoid pure, whereas $A/(f+2x_{1}\cdots x_{4})$ is perfectoid pure with $\operatorname{ppt}(A/f,p)=0$ Proposition 6.3(2) and (3).. Indeed, $f\in\mathfrak{m}^{[p]}$ and $\Delta(f)\in\mathfrak{m}^{[p^{2}]}$.

### v2

@@ 7. Examples
Remark 7.2.
Let $f\in A^{\prime}:=W(k)[x_{1},\ldots,x_{N}]$ be a homogeneous element. Then $A/f$ is perfectoid pure if and only if $(A^{\prime}/f)^{\wedge p}$ is perfectoid pure, and moreover we have
$\operatorname{ppt}(A/f,p)=\operatorname{ppt}((A^{\prime}/f)^{\wedge p},p)$
by Proposition 6.2 together with the proof of [p-pure]*Lemma 4.8.
@@
With the notation as in 7.1, let $f=x_{1}^{N}+\cdots+x_{N}^{N}$ and assume $p>N$. For each $e\geq 1$, take an integer $0\leq s_{e}\leq N-2$ satisfying $s_{e}+1\equiv p^{e}\pmod{N}$. Then the splitting–order sequence of $f$ is $\boldsymbol{s}(f)=(0,s_{1},s_{2},\ldots)$. In particular, $A/f$ is perfectoid pure.
@@
With the notation as in 7.1, assume that $f\in\mathfrak{m}^{[p]}$.
@@
    Then $A/f$ is not perfectoid pure, whereas $A/(f+2x_{1}\cdots x_{4})$ is perfectoid pure with $\operatorname{ppt}(A/f,p)=0$ Proposition 7.4 (2) and (3). Indeed, $f\in\mathfrak{m}^{[p]}$ and $\Delta(f)\in\mathfrak{m}^{[p^{2}]}$.
@@
Example 7.9.
Let $N=5$ and $f=x_{1}^{4}+\cdots+x_{5}^{4}$.
  • If $p=2$, then $f\in\mathfrak{m}^{[2]}$ and $\Delta(f)\in\mathfrak{m}^{[4]}$. Hence $R$ is not perfectoid pure by Proposition 7.4 (2).
  • If $p\equiv 1\pmod{4}$ or $p>7$, then $R$ is perfectoid pure with $\operatorname{ppt}(R,p)=1$. Indeed, $f^{p-1}$ contains either $(x_{1}\cdots x_{4})^{p-1}$ or $(x_{1}\cdots x_{4})^{p-3}x_{5}^{8}$, both of which are not contained in $\mathfrak{m}^{[p]}$.
  • If $p=3$, then $\boldsymbol{s}(f)=(0,2,0,0,\ldots)$. In particular, $R$ is perfectoid pure with $\operatorname{ppt}(R,p)=1/3$. Since $u(F_{*}f^{p-1})=\mathfrak{m}^{2}$, we have $m=1$ and t=2t=-2, where $m,t$ are defined in Theorem 6.4. Hence it suffices to show that $s_{1}=2$ and $s_{2}=0$ by Theorem 6.4. Since $f\in\mathfrak{m}^{[3]}$, we have $s_{1}=2$. Moreover,
    $p^{2}f^{p^{2}-2p-1}\Delta(f)^{2}$ $=f^{p^{2}-2p-1}\bigl(f^{p}-(x_{1}^{4p}+\cdots+x_{5}^{4p})\bigr)^{2}$
    $\equiv f^{p^{2}-1}\pmod{(x_{1}^{9},\ldots,x_{5}^{9})}$.
    The $p$-adic order of the coefficient of $(x_{1}\cdots x_{4})^{p^{2}-1}$ in $f^{p^{2}-1}$ is $2$; hence $f^{p^{2}-p-1}\Delta(f)$ is not contained in $\mathfrak{m}^{[p^{2}]}$. Thus $I(2,1)\nsubseteq\mathfrak{m}^{[p^{2}]}$, as desired.
  • If $p=7$, then $\boldsymbol{s}(f)=(0,1,0,0,\ldots)$. In particular, $R$ is perfectoid pure with $\operatorname{ppt}(R,p)=6/7$. Since $u(F_{*}f^{p-1})=\mathfrak{m}$, we have $m=1$ and t=1t=-1, where $m,t$ are defined in Theorem 6.4. Hence it suffices to show that $s_{1}=1$ and $s_{2}=0$ by Theorem 6.4. Since $f^{p-1}\in\mathfrak{m}^{[p]}$ and
    $f^{p-2}\equiv(x_{1}\cdots x_{5})^{4}\notin\mathfrak{m}^{[p]}\pmod{\mathfrak{m}^{[p]}}$,
    we obtain $s_{1}=1$. By the same argument as in (3), it suffices to show that the $p$-adic order of the coefficient of $x_{1}^{48}x_{2}^{48}x_{3}^{40}x_{4}^{28}x_{5}^{28}$ in $f^{p^{2}-1}$ is $1$. Indeed, this coefficient is
    $\frac{(p^{2}-1)!}{12!\,12!\,10!\,7!\,7!},$
    whose $p$-adic order is $1$, as required.
In conclusion, if $p>2$, then $R$ is perfectoid BCM-regular by Theorem 6.4.